### v1

# Effective Field Neural Network

Abstract
In recent years, with the rapid development of machine learning, physicists have been exploring its new applications in solving or alleviating the curse of dimensionality in many-body problems. In order to accurately reflect the underlying physics of the problem, domain knowledge must be encoded into the machine learning algorithms. In this work, inspired by field theory, we propose a new set of machine learning models called effective field neural networks (EFNNs) that can automatically and efficiently capture important many-body interactions through multiple self-refining processes. Taking the classical $3$-spin infinite-range model and the quantum double exchange model as case studies, we explicitly demonstrate that EFNNs significantly outperform fully-connected deep neural networks (DNNs) and the effective model. Furthermore, with the help of convolution operations, the EFNNs learned in a small system can be seamlessly used in a larger system without additional training and the relative errors even decrease, which further demonstrates the efficacy of EFNNs in representing core physical behaviors.
The collective behaviors of many interacting particles—such as spins, molecules, and atoms—give rise to some of the most intriguing phenomena in condensed matter physics. However, theoretical and computational studies of many-body problems often confront the curse of dimensionality and one must devise strategies to circumvent it. Machine learning has emerged as a powerful tool for extracting effective features from high-dimensional datasets, garnering significant attention for its ability to tackle long-standing problems in condensed matter physics. Numerous machine learning techniques have been applied to a wide range of classical and quantum many-body systems, including classifying phases of matter with supervised [1, 2, 3] and unsupervised machine learning [4, 5, 6, 7, 8], representing many-body quantum states [9, 10, 11, 12, 13], accelerating Monte-Carlo simulations [14, 15, 16, 17], fitting high-dimensional potential energy surfaces [18, 19, 20, 21, 22] and searching for spin glass ground states [23].
Standard deep neural networks (DNNs) have been shown to be ineffective in solving many-body problems unless augmented with physical knowledge. For example, mapping the spin configuration to energy in an $8\times 8$ classical two-dimensional Ising model requires highly complex DNN structures [24]. Consequently, extensive research has focused on encoding physics into neural network architectures to enhance their effectiveness in physical applications, such as incorporating physical laws [25, 26, 27, 28] and physical constraints [29, 30], mimicking physical processes [31], accounting for symmetry [32] and integrating graph structure in atomic bonding [33, 34, 35, 36, 37, 38, 39], employing renormalization group procedures [40]. These physics-encoded neural networks have found applications in computing effective Hamiltonians [41, 42, 43], property prediction [41, 36, 43, 44, 45, 46, 47, 48, 49, 39], structure search [37, 38], etc. These methods, although often outperforming standard DNNs, tend to rely on intuitive and simplified imitations of physical principles, and still remain elusive in revealing the underlying the many-body interactions.
In this work, inspired by field theory, we propose a new class of machine learning models, called effective field neural networks (EFNNs). The core motivation is to decompose many-body interactions into single quasi-particle representations governed by an emergent effective field. Different from standard DNNs, where an adjacent layer relies only on the previous layer, EFNNs incorporate the initial feature values at every layer to form the self-similarity structure with parameters trains via a recursive self-refining process, in the spirit of renormalization group theory, and hence can effectively capture the underlying physics of many-body interactions. Unlike approaches such as FermiNet [50], which directly parameterizes high-dimensional wave-functions without explicit separation of quasi-particles and fields, our method does not presuppose a fixed ansatz for the wavefunction. Instead, the quasi-particle and effective field representations are learned through an recursive self-refining process, mirroring the renormalization group (RG) framework in field theory, while leveraging the expressive power of deep neural networks. This recursive approach progressively refines the accuracy of both the quasi-particle and effective field descriptions.
Figure 1: Energy evaluation of a 2D Ising model reformulated as a neural network. (a) Calculation of the effective field by summing the interacting spins. Each interacting spin is multiplied by its corresponding effective field to obtain independent spin values, and the total energy is determined by summing up these independent spins. (b) Neural network representation of the energy evaluation process. $\odot$ represents the element-wise multiplication.
We first illustrate the EFNN architecture by reformulating the classical 2D Ising model within the EFNN framework. The Hamiltonian of a classical 2D Ising model is defined as $H(S)=-J\sum_{\langle ij\rangle}s_{i}s_{j}$, where $\langle ij\rangle$ denotes a summation over pairs of nearest-neighbor sites, $s_{i}=\pm 1$ represents the spin at site $i$, $J$ is the interaction strength between two spins, and $S$ is the collection of all spins. First, we define the effective field on spin $s_{i}$ as $\phi_{i}(S)$, which is equal to the sum of the nearest neighbor spins of $s_{i}$ multiplied by $-\frac{J}{2}$. Then, we define the independent spin (quasi-particle) as $s_{i}\phi_{i}(S)$, hence the total energy reads $H(S)=\sum_{i}s_{i}\phi_{i}(S)$. As illustrated schematically in Fig. 1(a), the interacting spins are first mapped to an effective field, then the effective field combining with the interacting spins is mapped to independent spins. Finally, a summation is performed over the independent spins, and the total energy is obtained. We reformulate the computational procedure within a DNN-like structure, depicted in Fig. 1(b). Here, $S_{0}=S$ is the input layer, representing the interacting spins. The effective field layer $F_{1}$ and the quasi-particle layer $S_{1}$ constitute a single field-particle (FP) layer. The evaluation of $S_{1}$ relies on both the effective layer $F_{1}$ and the interacting spin layer $S_{0}$, therefore a connection is required from $S_{0}$ to $S_{1}$, as well as from $F_{1}$ to $S_{1}$, distinguishing this architecture from a standard DNN. Finally, a summation is performed over $S_{1}$ to obtain the energy $E$.
Figure 2: Performance of EFNNs on a classical 3-spin infinite range model. (a) Architecture of the EFNN. (b) Performance of EFNNs and DNNs on the test set. (c) Performance of EFNNs as the neuron number increases.
In the classical 2D Ising model, both the effective field and quasi-particles can be exactly calculated, which, however, becomes infeasible for more complex many-body interactions. Following the spirit of renormalization from field theory, we can recursively refine the evaluations of the effective field and the quasi-particles. As shown in Fig. 2(a), we can extend the FP layer number in Fig. 1(b) to obtain the deep effective field neural networks. The evaluation procedure is detailed as
$F_{i}$ $=f_{i-1}(S_{i-1}),\,S_{i}=g_{i}(S_{0})\odot F_{i},\,E=q(S_{n}),$
where the function $f_{i-1}$ maps the previous quasi-particle layer $S_{i-1}$ (for $i\geq 2$) or the initial layer $S_{0}$ (for $i=1$) to the effective field layer $F_{i}$. The input layer $S_{0}$ is first processed by $g_{i}$, then multiplied in elements by $F_{i}$ to produce the quasi-particle layer $S_{i}$. Each FP layer consists of $F_{i}$ and $S_{i}$. The final quasi-particle layer $S_{n}$ generates the output energy $E$ through the function $q$, which sums all elements of the last layer. Typically, $f_{i}$ and $g_{j}$ are nonlinear to enhance expressiveness, though linear functions may suffice for simpler models like the classical 2D Ising model with nearest-neighbor interactions.
We present two case studies to demonstrate the capability of EFNNs to capture many-body interactions. The first case study focuses on the classical 3-spin infinite range model in 1D, with Hamiltonian:
$H(S)$ $=-\sum_{i<j<k}J_{ijk}s_{i}s_{j}s_{k},$ (1)
where the spins take binary values $0,1$, $J_{ijk}$ are constants, and the system comprises 15 lattice sites. We train EFNNs to evaluate $H(S)$ in Eq. (1), and compare its performance with that of DNNs.
Figure 3: Computational workflow of EFNNs with symmetrization. (a) Symmetrization: Each of the three channels of $S_{0}$ ($S_{0,x}$, $S_{0,y}$, $S_{0,z}$) undergoes convolution and element-wise multiplication within the same channel. The resulting products are summed to form the symmetrization layer $T_{j}(S_{0})$. Subsequently, $T_{j}(S_{0})$ is passed through a sequence of layers—convolution, batch normalization, $\tanh$ activation, and another convolution—to produce $F_{1}$ for $j=0$ or $g_{j}(T_{j}(S_{0}))$ for $j=1,\ldots,n$. (b) Generation of quasi-particle layers: The transformed symmetrization layer $g_{j}(T_{j}(S_{0}))$ is element-wise multiplied with the effective field layer $F_{j}$, resulting in the quasi-particle layer $S_{j}$, $j=1,\ldots,n$. (c) Transformation to effective field layers: Each quasi-particle layer $S_{k-1}$ is processed through convolution, batch normalization, $\tanh$ activation, and another convolution to generate the effective field layer $F_{k}$, for $k=2,\ldots,n$. (d) Energy evaluation: The final quasi-particle layer $S_{n}$ is transformed into a single-channel matrix. An element-wise summation of this matrix yields the scalar energy prediction $E$. These images are generated after training the model with $C=10$ channels and $n=2$ layers. Only the first five channels are displayed for clarity, using the 576th sample from the test set as input $S_{0}$.
We generate 16000 training and 4000 test samples using Eq. (1). Utilizing the EFNN architecture in Fig. 2(a), the nonlinear mappings $f_{i}$ and $g_{j}$ are constructed by sequential linear layers with $\tanh$ activations. The network is trained with the Adam optimizer and a decaying learning rate. Performance is evaluated by the relative error, defined as the square root of the MSE on the test set divided by the absolute mean of $H(S)$. We compare EFNNs to standard DNNs, using 18 neurons per layer and varying the number of FP layers ($n=1,2,3$). With one FP layer, the DNN achieves approximately a $7\times 10^{-2}$ relative error, while the EFNN attains around $4\times 10^{-2}$. As the number of FP layers increases, the DNN’s error rapidly worsens, whereas the EFNN’s error decreases to about $1\times 10^{-2}$ with three FP layers (Fig. 2(b)). Additionally, Fig. 2(c) shows that increasing the number of neurons in the EFNN from 15 to 150 leads to an algebraic decrease in relative error, reaching a minimum around 150 neurons for all tested FP layers, and optimal performance is achieved with three FP layers. These results demonstrate that EFNNs effectively capture many-body interactions through multiple feed-forward paths, improving accuracy as layers and neurons increase. In contrast, DNNs lose their ability to fit the energy accurately with additional layers. Both the number of FP layers and the internal neuron count are crucial for the EFNNs’ high performance.
We proceed to our second, more challenging case study: evaluating the Monte-Carlo energy of the quantum double exchange model [51, 52, 53] at finite temperature. The Hamiltonian on an $N\times N$ lattice reads
$H(S)=-t\sum_{\langle i,j\rangle,\alpha}\left(\hat{c}_{i\alpha}^{\dagger}\hat{c}_{j\alpha}+\text{h.c.}\right)-\frac{J}{2}\sum_{i,\alpha,\beta}\vec{s}_{i}\cdot\hat{c}_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}\hat{c}_{i\beta},$ (2)
where ${\langle ij\rangle}$ denotes nearest neighbors, $\hat{c}_{i\alpha}$ is the fermion annihilation operator with spin ${\alpha}$ at site $i$, $\vec{\sigma}$ are the Pauli matrices, and $\vec{s}_{i}\in\mathbb{S}^{2}$ is a 3D classical unit vector representing the local spin at site $i$, the classical spin interacts with the fermions through an on-site coupling, and $S$ is the collection of all classical spins $\vec{s}_{i}$. We set $N=10$, $J=16t$, $\mu=-8.3t$, $T=0.1t$, $t=1$.
At finite temperature $T$, the Monte-Carlo energy is defined as $E_{\operatorname{MC}}(S)=-T\sum_{n}\log(1+e^{-\frac{1}{T}E_{n}(S)})$, where $\{E_{n}(S)\}$ are the eigen-energies of $H(S)$ in Eq. (2) for a given spin configuration $S$. This quantity proves to be useful in Monte-Carlo simulations of the double exchange model, as it can replace exact diagonalization in Determinant Quantum Monte Carlo (DQMC) computations [54]. Diagonalizing a $2N^{2}\times 2N^{2}$ Hermitian matrix for Eq. (2) has a computational complexity of $O(N^{6})$ for each spin update. Both effective models and neural networks can significantly reduce this complexity. We compare the performance of an effective model incorporating 4-body RKKY-type interactions [55, 56, 57] with that of EFNNs adopting symmetrization. The effective model is expressed as
$E_{\operatorname{eff}}(S)$ $\sim J_{0}+\sum_{ij}g_{ij}\vec{s}_{i}\cdot\vec{s}_{j}+\sum_{ijkl}g_{ijkl}(\vec{s}_{i}\cdot\vec{s}_{j})(\vec{s}_{k}\cdot\vec{s}_{l}).$ (3)
Here, the dot products include all pairs of spins on the lattice, including self-interactions. Clearly, the number of parameters factorially increases with $N$ and the effective model cannot be solved exactly for a large $N$. To simplify, we divide the $10\times 10$ lattice into four $5\times 5$ tiles, restricting two- and four-body interactions within each tile.
The spin interactions in Eq. (3) are exclusively expressed as dot products to satisfy the $\operatorname{O}(3)$ symmetry of Eq. (2). Inspired by this symmetry, we incorporate symmetrization layers into the EFNN architecture, depicted in Fig. 3(a). The symmetrization is achieved by summing dot products of each channel of $S_{0}$ after convolution:
$T_{j}(S_{0})$ $=\sum_{k=x,y,z}\operatorname{conv}(S_{0,k})\odot\operatorname{conv}(S_{0,k}),\,\,j=0,1,\ldots,n.$
Here, each $S_{0,k}\in\mathbb{R}^{N\times N}$ is convolved into $\mathbb{R}^{C\times N\times N}$. Supplementary Material [58] details the convolution process and proves that $T_{j}$ remains invariant under $\operatorname{O}(3)$ transformations. As shown in Fig. 4(a), in the first FP layer, $F_{1}$ and $S_{1}$ are initialized as
$F_{1}$ $=f_{0}(T_{0}(S_{0})),\,\,S_{1}=g_{1}(T_{1}(S_{0}))\odot F_{1}.$
For subsequent layers $j=2,\ldots,n$, the effective field and quasi-particle layers are
$F_{j}$ $=f_{j-1}(S_{j-1}),\,\,S_{j}=g_{j}(T_{j}(S_{0}))\odot F_{j}.$
The mappings $f_{j}$ and $g_{k}$ consist of a convolutional layer, batch normalization, a $\tanh$ activation, and another convolutional layer. Fig. 3 illustrates the computational workflow: (a) symmetrization, (b) generation of quasi-particle layers, (c) transformation to effective field layers for subsequent FP layers, and (d) evaluation of energy. For the energy function $q$, after passing through the aforementioned layers, an extra summation layer generates a scalar value as the predicted energy. Specifically, the final quasi-particle layer $S_{n}$ is transformed into a single-channel matrix, and an element-wise summation is performed to produce the scalar energy prediction $E$, as shown in Fig. 3(d). The layer dimensions are as follows: input layer $S_{0}\in\mathbb{R}^{3\times N\times N}$, intermediate layers $F_{j}$, $S_{j}\in\mathbb{R}^{C\times N\times N}$, for $j=1,\ldots,n$, and output $E=q(S_{n})\in\mathbb{R}$.
Figure 4: Relative error of EFNNs. (a) Architecture of EFNN incorporating symmetrization layers $T_{j}$, $j=0,1,\ldots,n$. (b) Comparison of EFNNs’ and the effective model’s performance on a test set with lattice size $N=10$, for channel numbers $C=10,15,20,25$. (c) Performance of the EFNN model, trained on a $10\times 10$ lattice with $C=20$ channels, when applied to larger lattice systems.
We generate a total of $2\times 10^{6}$ data points and split them into 80% for training and 20% for testing. These data are obtained by exact diagonalization of the Hamiltonian in Eq. (2). For the effective model in Eq. (3), we perform a linear regression to obtain the coefficients. The EFNNs with symmetrization are trained using networks with $C=10,15,20,25$ channels and $n=1,2,3$ FP layers. As shown in Fig. 4(b), the effective model achieves a relative error of approximately $2\times 10^{-2}$. In contrast, the EFNNs with symmetrization consistently reduce the relative error below $3\times 10^{-3}$ as $C$ increases from 10 to 25, for $n=1,2,3$. Notably, the EFNNs outperform the effective model by a factor of six while utilizing significantly fewer parameters, underscoring their capability to capture high-order interactions in quantum many-body systems. Their parameter count scales as $O(nC^{2})$, resulting in 26250, 49560, and 72870 parameters for $n=1,2,3$ at $C=10$. In stark contrast, the perturbative expansion in Eq. (3) scales factorially as $O\bigg{(}\binom{\frac{1}{2}N^{4}+\frac{1}{2}N^{2}+\frac{1}{2}k-1}{\frac{1}{2}k}\bigg{)}$, with $k$ denoting the interaction order—a scaling that can lead to divergence at higher orders. The full effective model in Eq. (3) contains 12758826 parameters, which reduces to 213201 when interactions are restricted within each tile. Furthermore, EFNNs effectively select the most significant interactions automatically, an essential aspect of renormalization. By renormalizing the interactions and prioritizing key terms, EFNNs avoid both divergence and explosive growth in the number of parameters.
We further assess the extrapolation performance of EFNNs with symmetrization with the model trained on a $10\times 10$ lattice and tested on systems of $N=15,20,25,30,35$, and 40. Using the relative error of energy as the performance metric, Fig. 4(c) illustrates that for $N=10$, all EFNNs accurately predict the energy, consistent with Fig. 4(b). As the lattice size increases, the relative error slightly decreases for all FP layers with $n=1,2,3$. This occurs because the absolute error increases only marginally with $N$, while the absolute value of the average energy grows significantly, thereby reducing the relative error. Moreover, EFNNs with $n=2$ and $n=3$ FP layers outperform those with a single FP layer, demonstrating enhanced renormalization capability with additional layers. Notably, even for larger systems, all models maintain relative errors below $3\times 10^{-3}$, with the $n=2$ configuration achieving a minimum relative error of $1\times 10^{-3}$. These results underscore the EFNN’s efficacy: the convolution operations with padding under boundary condition PBC effectively capture interactions between localized and neighboring spins, and the renormalization through multiple FP layers ensures robust energy predictions across varying lattice sizes.
Figure 5: Continued function representation of an EFNN with 3 FP layers. (a) In an EFNN, the initial layer $S_{0}$ is recursively integrated into every quasi-particle layer. After a mapping, $S_{0}$ is multiplied with subsequent layers, forming a continued function representation (see right-hand side equation). (b) A typical architecture of ResNet, only one skip connection starts from $S_{0}$ and summation is used, which limits its renormalization capability. (c) Removing the $S_{0}$ connections transforms an EFNN into a standard DNN, where $S_{0}$ only appears at the beginning of the iterations, resulting in significantly diminished expressive power.
We give some discussions and conclude this paper. In both case studies, we have demonstrated that increasing the number of layers and neurons (or channels in convolutional layers) enhances the neural network’s accuracy, aligning with well-established findings in fields like computer vision. EFNN’s self-similar structure mirrors the renormalization process—a non-perturbative method for evaluating complicated functions, such as the Yang-Mills $\beta$ function [59] or Becke’s correction to the LDA functional in DFT [60]. This approximation framework stems from Padé approximants [61] using rational functions, which can also be expressed as continued fractions—an important and specific example of continued functions [62], known for superior convergence with much fewer parameters [63, 64, 65, 66]. Actually, EFNNs possess exactly the structure of continued functions, as shown in Fig. 5(a). In an EFNN, the initial feature layer $S_{0}$ connects recursively to every quasi-particle layer via mappings $g_{j}$; each nonlinear mapping $f_{j-1}$ is multiplied with $g_{j}(S_{0})$. In addition, EFNNs employ the $\tanh$ activation function within their renormalization-like structure, making them compatible with standard deep learning frameworks and enabling training via stochastic gradient descent with back-propagation. This contrasts with Padé approximants, which typically use the inverse function and hence usually determine parameters through perturbative series matching due to potential singularity in the gradient descent. As shown in Fig. 5(b), ResNets  [67, 68, 69, 70] adopt a similar but fundamentally different skip-like connections from EFNNs. In a ResNet, $S_{0}$ is skip-connected to later layers only once, with later skip connections starting in the middle, and its summation-based aggregation fails to meet the continued function formulation—resulting in suboptimal performance to characterize the many-body interactions. Standard DNNs (Fig. 5(c)) lack skip connections entirely, and $S_{0}$ only appears in the beginning of iterations, resulting in a very limited expressive capability. Moreover, by partitioning intermediate layers into effective field and quasi-particle layers, EFNNs provide a clear physical interpretation: quasi-particles emerge from the dot product of interacting particles (via a mapping) and effective fields, which is absent in ResNets and standard DNNs.
In conclusion, inspired by field theory, we have proposed a novel deep neural network architecture called EFNN, which effectively captures high-order interactions in both classical and quantum models. EFNN’s high accuracy and efficiency arise from the recursive refinement of effective fields and the construction of quasi-particles that encapsulate complex many-body effects. As discussed earlier, the connections from the true particle layer $S_{0}$ to the quasi-particle layers (or the preceding symmetrization layers) establish a structure of continued functions, and have renormalization power analogous to Padé approximants. Once the effective fields and quasi-particles are properly renormalized, the trained model can accurately predict the energy for a system with complicated many-body interactions. Furthermore, EFNNs maintain high accuracy even when applied to significantly larger systems. In future work, we aim to integrate Padé approximation and continued function theories more rigorously into the EFNN framework and derive error estimates for the neural network’s predictions.
Acknowledgments—This work is supported by the Hong Kong Research Grants Council (16306220) and National Key R&\&&D Program of China (2021YFA1401500). Y.Z. is supported by the Max Planck Partner lab on quantum materials from Max Planck Institute Chemical Physics of Solids.

### v2

@@
-Effective Field Neural Network
+Renormalization-Inspired Effective Field Neural Networks for Scalable Modeling of Classical and Quantum Many-Body Systems
 Abstract
-In recent years, with the rapid development of machine learning, physicists have been exploring its new applications in solving or alleviating the curse of dimensionality in many-body problems.
-In order to accurately reflect the underlying physics of the problem, domain knowledge must be encoded into the machine learning algorithms.
-In this work, inspired by field theory, we propose a new set of machine learning models called effective field neural networks (EFNNs) that can automatically and efficiently capture important many-body interactions through multiple self-refining processes.
-Taking the classical $3$-spin infinite-range model and the quantum double exchange model as case studies, we explicitly demonstrate that EFNNs significantly outperform fully-connected deep neural networks (DNNs) and the effective model.
-Furthermore, with the help of convolution operations, the EFNNs learned in a small system can be seamlessly used in a larger system without additional training and the relative errors even decrease, which further demonstrates the efficacy of EFNNs in representing core physical behaviors.
+We introduce Effective Field Neural Networks (EFNNs), a new architecture based on continued functions—mathematical tools used in renormalization to handle divergent perturbative series.
+Our key insight is that neural networks can implement these continued functions directly, providing a principled approach to many-body interactions.
+Testing on three systems (a classical 3-spin infinite- range model, a continuous classical Heisenberg spin system, and a quantum double exchange model), we find that EFNN outperforms standard deep networks, ResNet, and DenseNet.
+Most striking is EFNN’s generalization: trained on 10×10 lattices, it accurately predicts behavior on systems up to 40×40 with no additional training—and the accuracy improves with system size, with a computational time speed-up of $10^{3}$ compared to ED for $40\times 40$ lattice.
+This demonstrates that EFNN captures the underlying physics rather than merely fitting data, making it valuable beyond many-body problems to any field where renormalization ideas apply.
 The collective behaviors of many interacting particles—such as spins, molecules, and atoms—give rise to some of the most intriguing phenomena in condensed matter physics.
 However, theoretical and computational studies of many-body problems often confront the curse of dimensionality and one must devise strategies to circumvent it.
-Machine learning has emerged as a powerful tool for extracting effective features from high-dimensional datasets, garnering significant attention for its ability to tackle long-standing problems in condensed matter physics.
-Numerous machine learning techniques have been applied to a wide range of classical and quantum many-body systems, including classifying phases of matter with supervised [1, 2, 3] and unsupervised machine learning [4, 5, 6, 7, 8], representing many-body quantum states [9, 10, 11, 12, 13], accelerating Monte-Carlo simulations [14, 15, 16, 17], fitting high-dimensional potential energy surfaces [18, 19, 20, 21, 22] and searching for spin glass ground states [23].
-Standard deep neural networks (DNNs) have been shown to be ineffective in solving many-body problems unless augmented with physical knowledge.
-For example, mapping the spin configuration to energy in an $8\times 8$ classical two-dimensional Ising model requires highly complex DNN structures [24].
-Consequently, extensive research has focused on encoding physics into neural network architectures to enhance their effectiveness in physical applications, such as incorporating physical laws [25, 26, 27, 28] and physical constraints [29, 30], mimicking physical processes [31], accounting for symmetry [32] and integrating graph structure in atomic bonding [33, 34, 35, 36, 37, 38, 39], employing renormalization group procedures [40].
+Machine learning has emerged as a powerful tool for extracting effective features from high-dimensional datasets and tackling long-standing problems in condensed matter physics.
+Numerous machine learning techniques have been applied to a wide range of classical and quantum many-body systems, including classifying phases of matter with supervised [1, 2, 3] and unsupervised machine learning [4, 5, 6, 7, 8], representing many-body quantum states [9, 10, 11, 12, 13], accelerating Monte Carlo simulations [14, 15, 16, 17], fitting high-dimensional potential energy surfaces [18, 19, 20, 21, 22] and searching for spin glass ground states [23].
+Standard deep neural networks (DNNs) are ineffective in solving many-body problems unless augmented with physical knowledge.
+For example, an 8×8 classical two-dimensional Ising model requires highly complex DNN structures  [24].
+Consequently, extensive research has focused on encoding physics into neural network architectures to enhance their effectiveness in physical applications, including incorporating physical laws and constraints [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
 These physics-encoded neural networks have found applications in computing effective Hamiltonians [41, 42, 43], property prediction [41, 36, 43, 44, 45, 46, 47, 48, 49, 39], structure search [37, 38], etc.
-These methods, although often outperforming standard DNNs, tend to rely on intuitive and simplified imitations of physical principles, and still remain elusive in revealing the underlying the many-body interactions.
-In this work, inspired by field theory, we propose a new class of machine learning models, called effective field neural networks (EFNNs).
+These methods, although often outperforming standard DNNs, tend to rely on intuitive and simplified imitations of physical principles, and still remain elusive in revealing the underlying many-body interactions.
+In this work, we propose Effective Field Neural Networks (EFNNs), a new architecture inspired by field theory.
 The core motivation is to decompose many-body interactions into single quasi-particle representations governed by an emergent effective field.
-Different from standard DNNs, where an adjacent layer relies only on the previous layer, EFNNs incorporate the initial feature values at every layer to form the self-similarity structure with parameters trains via a recursive self-refining process, in the spirit of renormalization group theory, and hence can effectively capture the underlying physics of many-body interactions.
+Unlike standard DNNs where adjacent layers rely only on previous layers, EFNNs incorporate initial features at every layer, forming a self-similar structure trained via recursive self-refining that captures many-body interactions through renormalization principles.
 Unlike approaches such as FermiNet [50], which directly parameterizes high-dimensional wave-functions without explicit separation of quasi-particles and fields, our method does not presuppose a fixed ansatz for the wavefunction.
-Instead, the quasi-particle and effective field representations are learned through an recursive self-refining process, mirroring the renormalization group (RG) framework in field theory, while leveraging the expressive power of deep neural networks.
+Instead, the quasi-particle and effective field representations are learned through a recursive self-refining process, mirroring the renormalization group (RG) framework in field theory, while leveraging the expressive power of deep neural networks.
 This recursive approach progressively refines the accuracy of both the quasi-particle and effective field descriptions.
@@
 $\odot$ represents the element-wise multiplication.
+Figure 2: Performance of EFNNs on a classical 3-spin infinite range model and classical Heisenberg spin system.
+(a) Architecture of the EFNN.
+(b) Performance of EFNNs, DenseNets, ResNets, DNNs on 3-spin infinite range model.
+(c) Performance of EFNNs, DenseNets, ResNets, DNNs on classical Heisenberg spin system.
 We first illustrate the EFNN architecture by reformulating the classical 2D Ising model within the EFNN framework.
-The Hamiltonian of a classical 2D Ising model is defined as $H(S)=-J\sum_{\langle ij\rangle}s_{i}s_{j}$, where $\langle ij\rangle$ denotes a summation over pairs of nearest-neighbor sites, $s_{i}=\pm 1$ represents the spin at site $i$, $J$ is the interaction strength between two spins, and $S$ is the collection of all spins.
-First, we define the effective field on spin $s_{i}$ as $\phi_{i}(S)$, which is equal to the sum of the nearest neighbor spins of $s_{i}$ multiplied by $-\frac{J}{2}$.
-Then, we define the independent spin (quasi-particle) as $s_{i}\phi_{i}(S)$, hence the total energy reads $H(S)=\sum_{i}s_{i}\phi_{i}(S)$.
+The Hamiltonian of a classical 2D Ising model is defined as $H(S)=-J\sum_{\langle i,j\rangle}s_{i}s_{j}$, where $\langle i,j\rangle$ denotes a summation over pairs of nearest-neighbor sites, $s_{i}=\pm 1$ represents the spin at site $i$, $J$ is the interaction strength between two spins, and $S$ is the collection of all spins.
+We define the effective field on spin $s_{i}$ as $\phi_{i}(S)=-\frac{J}{2}\times(\textrm{sum of nearest neighbors})$, and the quasi-particle as $s_{i}\phi_{i}(S)$, yielding total energy $H(S)=\sum_{i}s_{i}\phi_{i}(S)$.
 As illustrated schematically in Fig.
-1(a), the interacting spins are first mapped to an effective field, then the effective field combining with the interacting spins is mapped to independent spins.
+1(a), the interacting spins are first mapped to an effective field, then the effective field element-wise multiplied with the interacting spins is mapped to independent spins.
 Finally, a summation is performed over the independent spins, and the total energy is obtained.
-We reformulate the computational procedure within a DNN-like structure, depicted in Fig.
-1(b).
-Here, $S_{0}=S$ is the input layer, representing the interacting spins.
+We reformulate this as a neural network structure (Fig.
+1(b)), where $S_{0}=S$ represents the interacting spins.
 The effective field layer $F_{1}$ and the quasi-particle layer $S_{1}$ constitute a single field-particle (FP) layer.
@@
 Finally, a summation is performed over $S_{1}$ to obtain the energy $E$.
-Figure 2: Performance of EFNNs on a classical 3-spin infinite range model.
-(a) Architecture of the EFNN.
-(b) Performance of EFNNs and DNNs on the test set.
-(c) Performance of EFNNs as the neuron number increases.
-In the classical 2D Ising model, both the effective field and quasi-particles can be exactly calculated, which, however, becomes infeasible for more complex many-body interactions.
+In the classical 2D Ising model, both the effective field and quasi-particles can be exactly calculated, which, however, becomes infeasible for more complicated many-body interactions.
 Following the spirit of renormalization from field theory, we can recursively refine the evaluations of the effective field and the quasi-particles.
+This recursive structure, when combined with the multiplicative interactions and self-similar connectivity pattern, implements a continued function—a mathematical tool from renormalization theory for handling divergent perturbative series (see Supplementary Material [51] for the complete derivation).
 As shown in Fig.
@@
 where the function $f_{i-1}$ maps the previous quasi-particle layer $S_{i-1}$ (for $i\geq 2$) or the initial layer $S_{0}$ (for $i=1$) to the effective field layer $F_{i}$.
-The input layer $S_{0}$ is first processed by $g_{i}$, then multiplied in elements by $F_{i}$ to produce the quasi-particle layer $S_{i}$.
+The input layer $S_{0}$ is first processed by $g_{i}$, then element-wise multiplied by $F_{i}$ to produce the quasi-particle layer $S_{i}$.
 Each FP layer consists of $F_{i}$ and $S_{i}$.
 The final quasi-particle layer $S_{n}$ generates the output energy $E$ through the function $q$, which sums all elements of the last layer.
-Typically, $f_{i}$ and $g_{j}$ are nonlinear to enhance expressiveness, though linear functions may suffice for simpler models like the classical 2D Ising model with nearest-neighbor interactions.
-We present two case studies to demonstrate the capability of EFNNs to capture many-body interactions. The first case study focuses on the classical 3-spin infinite range model in 1D, with Hamiltonian:
+The functions $f_{i}$ are nonlinear (specifically, we use hyperbolic tangent functions) to implement renormalization by ensuring finite values, while $g_{j}$ are nonlinear to enhance expressiveness.
+Linear functions may suffice only for simpler models like the classical 2D Ising model with nearest-neighbor interactions.
+We present three case studies to demonstrate the capability of EFNNs to capture many-body interactions.
+The first case study focuses on the classical 3-spin infinite-range model in 1D, with Hamiltonian:
 $H(S)$ $=-\sum_{i<j<k}J_{ijk}s_{i}s_{j}s_{k},$ (1)
-where the spins take binary values $0,1$, $J_{ijk}$ are constants, and the system comprises 15 lattice sites.
+where the spins take binary values $\{0,1\}$, $J_{ijk}$ are constants, and the system comprises 15 lattice sites.
 We train EFNNs to evaluate $H(S)$ in Eq.
-(1), and compare its performance with that of DNNs.
-Figure 3: Computational workflow of EFNNs with symmetrization.
-(a) Symmetrization: Each of the three channels of $S_{0}$ ($S_{0,x}$, $S_{0,y}$, $S_{0,z}$) undergoes convolution and element-wise multiplication within the same channel.
+(1), and compare its performance with those of DenseNets, ResNets, and DNNs.
+Figure 3: Symmetrization: Each of the three channels of $S_{0}$ ($S_{0,x}$, $S_{0,y}$, $S_{0,z}$) undergoes convolution and element-wise multiplication within the same channel.
 The resulting products are summed to form the symmetrization layer $T_{j}(S_{0})$.
-Subsequently, $T_{j}(S_{0})$ is passed through a sequence of layers—convolution, batch normalization, $\tanh$ activation, and another convolution—to produce $F_{1}$ for $j=0$ or $g_{j}(T_{j}(S_{0}))$ for $j=1,\ldots,n$.
-(b) Generation of quasi-particle layers: The transformed symmetrization layer $g_{j}(T_{j}(S_{0}))$ is element-wise multiplied with the effective field layer $F_{j}$, resulting in the quasi-particle layer $S_{j}$, $j=1,\ldots,n$.
-(c) Transformation to effective field layers: Each quasi-particle layer $S_{k-1}$ is processed through convolution, batch normalization, $\tanh$ activation, and another convolution to generate the effective field layer $F_{k}$, for $k=2,\ldots,n$.
-(d) Energy evaluation: The final quasi-particle layer $S_{n}$ is transformed into a single-channel matrix.
-An element-wise summation of this matrix yields the scalar energy prediction $E$.
-These images are generated after training the model with $C=10$ channels and $n=2$ layers.
-Only the first five channels are displayed for clarity, using the 576th sample from the test set as input $S_{0}$.
+Subsequently, $T_{j}(S_{0})$ is passed through a sequence of layers—convolution, $\tanh$ activation, and another convolution
+—to produce $F_{1}$ for $j=0$ or $g_{j}(T_{j}(S_{0}))$ for $j=1,\ldots,n$.
 We generate 16000 training and 4000 test samples using Eq.
 (1).
-Utilizing the EFNN architecture in Fig.
-2(a), the nonlinear mappings $f_{i}$ and $g_{j}$ are constructed by sequential linear layers with $\tanh$ activations.
+The nonlinear mappings $f_{i}$ and $g_{j}$ use sequential linear layers with tanh activations.
 The network is trained with the Adam optimizer and a decaying learning rate.
-Performance is evaluated by the relative error, defined as the square root of the MSE on the test set divided by the absolute mean of $H(S)$.
-We compare EFNNs to standard DNNs, using 18 neurons per layer and varying the number of FP layers ($n=1,2,3$).
-With one FP layer, the DNN achieves approximately a $7\times 10^{-2}$ relative error, while the EFNN attains around $4\times 10^{-2}$.
-As the number of FP layers increases, the DNN’s error rapidly worsens, whereas the EFNN’s error decreases to about $1\times 10^{-2}$ with three FP layers (Fig.
-2(b)).
-Additionally, Fig.
-2(c) shows that increasing the number of neurons in the EFNN from 15 to 150 leads to an algebraic decrease in relative error, reaching a minimum around 150 neurons for all tested FP layers, and optimal performance is achieved with three FP layers.
-These results demonstrate that EFNNs effectively capture many-body interactions through multiple feed-forward paths, improving accuracy as layers and neurons increase.
-In contrast, DNNs lose their ability to fit the energy accurately with additional layers.
-Both the number of FP layers and the internal neuron count are crucial for the EFNNs’ high performance.
-We proceed to our second, more challenging case study: evaluating the Monte-Carlo energy of the quantum double exchange model [51, 52, 53] at finite temperature.
-The Hamiltonian on an $N\times N$ lattice reads
+Performance is evaluated by the relative error, defined as the root mean square error (RMSE) on the test set divided by the absolute value of the mean of $H(S)$.
+We vary the number of network layers from 1 to 3 and the number of neurons from 15 to 150.
+The performance of EFNNs, DenseNets, ResNets, and DNNs is presented in Fig.
+2(b).
+The relative error decreases with more neurons for all networks, but EFNNs significantly outperform others as layers increase, with the relative error decreasing below $5\times 10^{-3}$ for EFNNs with 2 and 3 layers when the number of neurons $\geq 105$.
+The only exception is the ResNet with 3 layers and 150 neurons.
+In most cases, DenseNets, ResNets, and DNNs perform worse than 2-layer EFNNs.
+To test the capability of EFNNs beyond discrete systems, our second case study examines a classical Heisenberg spin system, using the same three-body interaction model in Eq.
+(1) with continuous spins, where $s_{i}$ takes continuous values in [1,1][-1,1] rather than discrete values 0,1{0,1}.
+We use the RMSE of the total energy as the performance metric, with results shown in Fig.
+2(c), since the average value of Eq.
+(1) is 0.
+Compared to the discrete case, the configuration space is now continuous, yet EFNNs with 2 and 3 layers still outperform all other networks.
+Fig.
+2(c) also shows the general trend that the error decreases with increasing number of neurons.
+The most accurate result is achieved by the 3-layer EFNN with 150 neurons, where the error of the total energy reduces to $4\times 10^{-2}$.
+These results demonstrate that EFNNs effectively capture many-body interactions through the renormalization procedure of their recursive continued-function structure; failing to incorporate this structure leads to inaccurate results.
+We show in the Supplementary Material [51] that DNNs completely lack the structure of a continued function, while DenseNets and ResNets misrepresent a continued function.
+We proceed to our third, more challenging case study: evaluating the Monte Carlo energy of the quantum double exchange model [52, 53, 54] at finite temperature.
+The Hamiltonian of an $N\times N$ lattice reads
 $H(S)=-t\sum_{\langle i,j\rangle,\alpha}\left(\hat{c}_{i\alpha}^{\dagger}\hat{c}_{j\alpha}+\text{h.c.}\right)-\frac{J}{2}\sum_{i,\alpha,\beta}\vec{s}_{i}\cdot\hat{c}_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}\hat{c}_{i\beta},$ (2)
-where ${\langle ij\rangle}$ denotes nearest neighbors, $\hat{c}_{i\alpha}$ is the fermion annihilation operator with spin ${\alpha}$ at site $i$, $\vec{\sigma}$ are the Pauli matrices, and $\vec{s}_{i}\in\mathbb{S}^{2}$ is a 3D classical unit vector representing the local spin at site $i$, the classical spin interacts with the fermions through an on-site coupling, and $S$ is the collection of all classical spins $\vec{s}_{i}$.
+where $\langle i,j\rangle$} denotes nearest neighbors, $\hat{c}_{i\alpha}$ is the electron annihilation operator with spin $\alpha$} at site $i$, $\vec{\sigma}$ are the Pauli matrices, and $\vec{s}_{i}\in\mathbb{S}^{2}$ is a 3D classical unit vector representing the local spin at site $i$, the classical spin interacts with the electrons through an on-site coupling, and $S$ is the collection of all classical spins $\vec{s}_{i}$.
 We set $N=10$, $J=16t$, $\mu=-8.3t$, $T=0.1t$, $t=1$.
-At finite temperature $T$, the Monte-Carlo energy is defined as $E_{\operatorname{MC}}(S)=-T\sum_{n}\log(1+e^{-\frac{1}{T}E_{n}(S)})$, where $\{E_{n}(S)\}$ are the eigen-energies of $H(S)$ in Eq.
-(2) for a given spin configuration $S$.
-This quantity proves to be useful in Monte-Carlo simulations of the double exchange model, as it can replace exact diagonalization in Determinant Quantum Monte Carlo (DQMC) computations [54].
-Diagonalizing a $2N^{2}\times 2N^{2}$ Hermitian matrix for Eq.
-(2) has a computational complexity of $O(N^{6})$ for each spin update.
-Both effective models and neural networks can significantly reduce this complexity.
-We compare the performance of an effective model incorporating 4-body RKKY-type interactions [55, 56, 57] with that of EFNNs adopting symmetrization.
-The effective model is expressed as
+At finite temperature $T$, the Monte Carlo energy $E_{\text{MC}}(S)=-T\sum_{n}\log(1+e^{-E_{n}(S)/T})$ determines the acceptance ratio in Monte Carlo simulations, where $E_{n}(S)$ are the eigenvalues of $H(S)$ with chemical potential in Eq.
+(2).
+Computing the exact value of $E_{\text{MC}}(S)$ requires diagonalization of Eq.
+(2) at each Monte Carlo update step, incurring $O(N^{6})$ computational complexity per step.
+With millions of Monte Carlo steps required for convergence, this approach becomes computationally prohibitive.
+An intuitive solution is to approximate $E_{\text{MC}}(S)$ using a perturbative expansion, yielding an effective model with RKKY-type interactions [55, 56, 57].
+This effective model with 2-body interactions has proven sufficiently accurate to replace exact diagonalization in Determinant Quantum Monte Carlo (DQMC) simulations [58].
+The effective model including up to 4-body interactions takes the form
 $E_{\operatorname{eff}}(S)$ $\sim J_{0}+\sum_{ij}g_{ij}\vec{s}_{i}\cdot\vec{s}_{j}+\sum_{ijkl}g_{ijkl}(\vec{s}_{i}\cdot\vec{s}_{j})(\vec{s}_{k}\cdot\vec{s}_{l}).$ (3)
 Here, the dot products include all pairs of spins on the lattice, including self-interactions.
-Clearly, the number of parameters factorially increases with $N$ and the effective model cannot be solved exactly for a large $N$.
-To simplify, we divide the $10\times 10$ lattice into four $5\times 5$ tiles, restricting two- and four-body interactions within each tile.
+The perturbative approach faces a fundamental dilemma: truncating at low orders fails to capture the full physics, while including higher-order terms leads to numerical divergence due to the factorial growth in the number of terms.
+EFNN addresses this challenge through renormalization.
 The spin interactions in Eq.
@@
 Inspired by this symmetry, we incorporate symmetrization layers into the EFNN architecture, depicted in Fig.
-3(a).
+3.
 The symmetrization is achieved by summing dot products of each channel of $S_{0}$ after convolution:
@@
 $F_{j}$ $=f_{j-1}(S_{j-1}),\,\,S_{j}=g_{j}(T_{j}(S_{0}))\odot F_{j}.$
-The mappings $f_{j}$ and $g_{k}$ consist of a convolutional layer, batch normalization, a $\tanh$ activation, and another convolutional layer.
-Fig.
-3 illustrates the computational workflow: (a) symmetrization, (b) generation of quasi-particle layers, (c) transformation to effective field layers for subsequent FP layers, and (d) evaluation of energy.
-For the energy function $q$, after passing through the aforementioned layers, an extra summation layer generates a scalar value as the predicted energy.
-Specifically, the final quasi-particle layer $S_{n}$ is transformed into a single-channel matrix, and an element-wise summation is performed to produce the scalar energy prediction $E$, as shown in Fig.
-3(d).
+The mappings $f_{j}$ and $g_{k}$ consist of a convolutional layer, a $\tanh$ activation, and another convolutional layer.
+For the mapping $q(S_{n})=E$, the final quasi-particle layer $S_{n}$ is convolved into a single-channel matrix, and an element-wise summation is performed to produce the scalar energy prediction $E$, see Supplementary Material [51].
 The layer dimensions are as follows: input layer $S_{0}\in\mathbb{R}^{3\times N\times N}$, intermediate layers $F_{j}$, $S_{j}\in\mathbb{R}^{C\times N\times N}$, for $j=1,\ldots,n$, and output $E=q(S_{n})\in\mathbb{R}$.
+For comparison, we also incorporate symmetrization layers and final summation $q$ into DenseNet, ResNet, and DNN architectures, detailed in Supplementary Material [51].
 Figure 4: Relative error of EFNNs.
 (a) Architecture of EFNN incorporating symmetrization layers $T_{j}$, $j=0,1,\ldots,n$.
-(b) Comparison of EFNNs’ and the effective model’s performance on a test set with lattice size $N=10$, for channel numbers $C=10,15,20,25$.
-(c) Performance of the EFNN model, trained on a $10\times 10$ lattice with $C=20$ channels, when applied to larger lattice systems.
+(b) Performance of EFNNs, DenseNets, ResNets, DNNs and effective model on a test set with lattice size $N=10$, for channel numbers $C=1,3,5,7,9$.
+(c) Generalization performance of networks trained on $10\times 10$ lattices ($C=3$) when applied to larger systems.
 We generate a total of $2\times 10^{6}$ data points and split them into 80% for training and 20% for testing.
 These data are obtained by exact diagonalization of the Hamiltonian in Eq.
-(2).
+(2) with chemical potential.
 For the effective model in Eq.
-(3), we perform a linear regression to obtain the coefficients.
-The EFNNs with symmetrization are trained using networks with $C=10,15,20,25$ channels and $n=1,2,3$ FP layers.
+(3), we perform linear regression to obtain the coefficients.
+While the full effective model would require 12758826 parameters, we reduce this to 213201 by restricting interactions within each $5\times 5$ tile.
+The EFNNs, DenseNets, ResNets, and DNNs with symmetrization are trained using networks with $C=1,3,5,7,9$ channels and $n=1,2,3$ layers, with parameter counts ranging from 200 to 30000.
 As shown in Fig.
-4(b), the effective model achieves a relative error of approximately $2\times 10^{-2}$.
-In contrast, the EFNNs with symmetrization consistently reduce the relative error below $3\times 10^{-3}$ as $C$ increases from 10 to 25, for $n=1,2,3$.
-Notably, the EFNNs outperform the effective model by a factor of six while utilizing significantly fewer parameters, underscoring their capability to capture high-order interactions in quantum many-body systems.
-Their parameter count scales as $O(nC^{2})$, resulting in 26250, 49560, and 72870 parameters for $n=1,2,3$ at $C=10$.
-In stark contrast, the perturbative expansion in Eq.
-(3) scales factorially as $O\bigg{(}\binom{\frac{1}{2}N^{4}+\frac{1}{2}N^{2}+\frac{1}{2}k-1}{\frac{1}{2}k}\bigg{)}$, with $k$ denoting the interaction order—a scaling that can lead to divergence at higher orders.
-The full effective model in Eq.
-(3) contains 12758826 parameters, which reduces to 213201 when interactions are restricted within each tile.
-Furthermore, EFNNs effectively select the most significant interactions automatically, an essential aspect of renormalization.
-By renormalizing the interactions and prioritizing key terms, EFNNs avoid both divergence and explosive growth in the number of parameters.
-We further assess the extrapolation performance of EFNNs with symmetrization with the model trained on a $10\times 10$ lattice and tested on systems of $N=15,20,25,30,35$, and 40.
-Using the relative error of energy as the performance metric, Fig.
-4(c) illustrates that for $N=10$, all EFNNs accurately predict the energy, consistent with Fig.
-4(b).
-As the lattice size increases, the relative error slightly decreases for all FP layers with $n=1,2,3$.
-This occurs because the absolute error increases only marginally with $N$, while the absolute value of the average energy grows significantly, thereby reducing the relative error.
-Moreover, EFNNs with $n=2$ and $n=3$ FP layers outperform those with a single FP layer, demonstrating enhanced renormalization capability with additional layers.
-Notably, even for larger systems, all models maintain relative errors below $3\times 10^{-3}$, with the $n=2$ configuration achieving a minimum relative error of $1\times 10^{-3}$.
-These results underscore the EFNN’s efficacy: the convolution operations with padding under boundary condition PBC effectively capture interactions between localized and neighboring spins, and the renormalization through multiple FP layers ensures robust energy predictions across varying lattice sizes.
-Figure 5: Continued function representation of an EFNN with 3 FP layers.
+4(b), the effective model achieves a relative error of approximately $2\times 10^{-2}$, whereas the neural networks achieve relative errors on the order of $10^{-3}$—approximately six times better using significantly fewer parameters.
+Among the neural networks, EFNNs with 1, 2, and 3 layers consistently outperform all other architectures, with some overlap observed between 1-layer EFNNs and 3-layer DNNs, as well as 1-, 2-, and 3-layer DenseNets at $C=1,3$.
+For EFNNs, the error systematically decreases as $C$ increases and as the number of layers increases from 1 to 3.
+This superior performance with fewer parameters shows that the EFNNs possess inherent renormalization capabilities beyond simple parameter fitting.
+We next test the generalization capability of these neural networks trained on $10\times 10$ lattices by evaluating their performance on systems up to $40\times 40$.
+We select networks with $C=3$ channels as a representative configuration and evaluate EFNN, DenseNet, ResNet, and DNN architectures across this range of lattice sizes.
+The convolution operations with padding under periodic boundary conditions (PBC) effectively capture interactions between localized and neighboring spins.
+As shown in Fig.
+4(c), all neural network architectures show improved performance on larger lattices, with relative errors decreasing as the system size increases from $10\times 10$ to $40\times 40$.
+The EFNNs exhibit the best extrapolation capability and renormalization ability: the 3-layer EFNN achieves the lowest errors across all lattice sizes, reaching a remarkable relative error of approximately $4\times 10^{-4}$ at $N=40$.
+The 2-layer EFNN mostly outperforms the 1-layer EFNN, while among non-EFNN architectures, only the 3-layer DNN has comparable performance to the 1-layer EFNN.
+All other configurations consistently produce higher errors, confirming EFNN’s superior ability to capture scale-invariant features.
+Notably, as demonstrated in Supplementary Material [51], for $N=40$, EFNN achieves $10^{3}$ speed-up compared to ED.
+Figure 5: Comparison of neural network structures.
 (a) In an EFNN, the initial layer $S_{0}$ is recursively integrated into every quasi-particle layer.
-After a mapping, $S_{0}$ is multiplied with subsequent layers, forming a continued function representation (see right-hand side equation).
-(b) A typical architecture of ResNet, only one skip connection starts from $S_{0}$ and summation is used, which limits its renormalization capability.
-(c) Removing the $S_{0}$ connections transforms an EFNN into a standard DNN, where $S_{0}$ only appears at the beginning of the iterations, resulting in significantly diminished expressive power.
+After a mapping, $S_{0}$ is multiplied with subsequent layers, forming a continued-function representation (see the equation above the figure).
+(b) A typical architecture of ResNet, only one skip connection starts from $S_{0}$ and summation is used, this is not a correct continued-function structure.
+(c) Removing the $S_{0}$ connections transforms an EFNN into a standard DNN, with no continued-function structure.
+(d) A typical architecture of DenseNet, $C$ is concatenation operation, this is not a correct continued-function structure.
 We give some discussions and conclude this paper.
-In both case studies, we have demonstrated that increasing the number of layers and neurons (or channels in convolutional layers) enhances the neural network’s accuracy, aligning with well-established findings in fields like computer vision.
-EFNN’s self-similar structure mirrors the renormalization process—a non-perturbative method for evaluating complicated functions, such as the Yang-Mills $\beta$ function [59] or Becke’s correction to the LDA functional in DFT [60].
-This approximation framework stems from Padé approximants [61] using rational functions, which can also be expressed as continued fractions—an important and specific example of continued functions [62], known for superior convergence with much fewer parameters [63, 64, 65, 66].
-Actually, EFNNs possess exactly the structure of continued functions, as shown in Fig.
-5(a).
-In an EFNN, the initial feature layer $S_{0}$ connects recursively to every quasi-particle layer via mappings $g_{j}$; each nonlinear mapping $f_{j-1}$ is multiplied with $g_{j}(S_{0})$.
-In addition, EFNNs employ the $\tanh$ activation function within their renormalization-like structure, making them compatible with standard deep learning frameworks and enabling training via stochastic gradient descent with back-propagation.
-This contrasts with Padé approximants, which typically use the inverse function and hence usually determine parameters through perturbative series matching due to potential singularity in the gradient descent.
-As shown in Fig.
-5(b), ResNets  [67, 68, 69, 70] adopt a similar but fundamentally different skip-like connections from EFNNs.
-In a ResNet, $S_{0}$ is skip-connected to later layers only once, with later skip connections starting in the middle, and its summation-based aggregation fails to meet the continued function formulation—resulting in suboptimal performance to characterize the many-body interactions.
-Standard DNNs (Fig.
-5(c)) lack skip connections entirely, and $S_{0}$ only appears in the beginning of iterations, resulting in a very limited expressive capability.
-Moreover, by partitioning intermediate layers into effective field and quasi-particle layers, EFNNs provide a clear physical interpretation: quasi-particles emerge from the dot product of interacting particles (via a mapping) and effective fields, which is absent in ResNets and standard DNNs.
-In conclusion, inspired by field theory, we have proposed a novel deep neural network architecture called EFNN, which effectively captures high-order interactions in both classical and quantum models.
-EFNN’s high accuracy and efficiency arise from the recursive refinement of effective fields and the construction of quasi-particles that encapsulate complex many-body effects.
-As discussed earlier, the connections from the true particle layer $S_{0}$ to the quasi-particle layers (or the preceding symmetrization layers) establish a structure of continued functions, and have renormalization power analogous to Padé approximants.
-Once the effective fields and quasi-particles are properly renormalized, the trained model can accurately predict the energy for a system with complicated many-body interactions.
-Furthermore, EFNNs maintain high accuracy even when applied to significantly larger systems.
-In future work, we aim to integrate Padé approximation and continued function theories more rigorously into the EFNN framework and derive error estimates for the neural network’s predictions.
-Acknowledgments—This work is supported by the Hong Kong Research Grants Council (16306220) and National Key R&\&&D Program of China (2021YFA1401500).
+In all three case studies—the classical 3-spin infinite-range model, continuous classical Heisenberg system, and quantum double exchange model—EFNNs consistently outperform DenseNets, ResNets, DNNs, and the effective model, achieving the lowest errors.
+Particularly striking is EFNN’s ability to generalize: trained on $10\times 10$ lattices, it accurately predicted energies for systems as large as $40\times 40$, while DenseNets, ResNets, and standard DNNs struggle with this extrapolation task.
+This superior performance isn’t accidental—it stems from EFNN’s mathematical foundation in continued functions.
+While conventional architectures like DenseNet and ResNet were designed through trial and error with various architectural innovations, EFNN’s structure directly implements continued functions, a mathematical framework that naturally captures many-body interactions up to infinite order.
+More importantly, for a perturbative expansion like Eq.
+(3), EFNN selects the most relevant terms via training, and expresses these terms in nonlinear continued function structure.
+Fig.
+5 compares these structures, demonstrating that only EFNN (Fig.
+5(a)) possesses the correct continued function structure.
+The Supplementary Material [51] provides a detailed mathematical explanation of why EFNN implements continued functions while ResNet and DenseNet do not.
+A special case of continued functions is the Padé approximant, which takes the form of a continued fraction and is ubiquitously used throughout physics and numerical methods for its rapid convergence, strong extrapolation capability, and compact representation [59].
+Continued functions generalize continued fractions by replacing the inversion operation with generic nonlinear functions.
+Here, we use tanh activation functions instead of inverses, which makes EFNN trainable through standard backpropagation—something impossible with traditional Padé approximants due to gradient singularities.
+This simple substitution preserves the mathematical power of continued functions while enabling practical neural network training.
+Notably, EFNN actually implements self-similar approximation [60], a powerful mathematical framework that has been successfully applied across diverse fields including high energy physics [61, 62, 63, 64, 65], condensed matter physics [66, 67], quantum mechanics [68, 69, 64, 70, 71, 72, 60], quantum field theory [73, 74, 60, 75], and applied mathematics [76, 66, 77, 78, 79, 80, 81, 82, 83, 84, 85].
+Previously, applying self-similar approximation required tedious manual calculations to obtain order by order matching; EFNN changes this by automating the process through neural network optimization.
+EFNN’s self-similar structure also naturally comes from renormalization group equation, as detailed in the Supplementary Material [51].
+In conclusion, we have developed EFNN, a neural network architecture inspired by field theory that captures high-order interactions in both classical and quantum models.
+The key to EFNN’s success lies in how it recursively refines effective fields and constructs quasi-particles to encode complex many-body effects.
+This partitioning into field and quasi-particle layers provides clear physical interpretation while maintaining computational efficiency.
+As we showed, the connections from the true particle layer $S_{0}$ to the quasi-particle layers create a continued function structure with renormalization properties similar to Padé approximants.
+What EFNN actually implements is self-similar approximation—a powerful mathematical technique that has been widely used in physics but previously required manual calculations for each problem.
+Once properly trained, this architecture learns effective representations that accurately predict energies even for systems with complicated many-body interactions.
+Our tests on three diverse systems—the classical 3-spin infinite-range model, continuous classical Heisenberg spin system, and quantum double exchange model—demonstrate that EFNN consistently outperforms DenseNet, ResNet, and standard DNNs.
+Most remarkably, EFNNs trained on small systems improve accuracy on larger ones, opening possibilities for training on expensive DFT data from small lattices then inferring on much larger systems.
+Looking ahead, we plan to develop theoretical error bounds and apply EFNN to other challenging problems in physics.
+DATA AVAILABILITY
+The data that support the findings of this study are openly available in Zenodo at Ref. [86].
+Acknowledgments—This work is supported by the Hong Kong Research Grants Council (16306220) and National Key R&\&D Program of China (2021YFA1401500).
 Y.Z.